(x+6)^2=3(x+12)-x^2
Answer by RAY100(1637)

(x+6)^2=3(x+12)-x^2
x^2+12x+36 =3x+36-x^2
2x^2+9x=0
x(2x+9)=0
x= 0, (-9/2) ANSWER
check
x=0
6^2=3*12
36=36 ok
x=(-9/2)
(-4.5+6)^2=3(-4.5+12)-(-4.5)^2
1.5^2=3*7.5-20.25
2.25=22.5-20.25=2.25 ok